OpenStudy (anonymous):

write down the expansion of sin(A+B) and hence find sin 15 deg in simplest surd form

OpenStudy (vishweshshrimali5):

sin(A+B) = sinAcosB + cosAsinB

OpenStudy (vishweshshrimali5):

sin(15) = sin(45-30) use sin(A-B) = sin A cos B - sin B cos A put A = 45 and B = 30

OpenStudy (vishweshshrimali5):

Got that ?

OpenStudy (anonymous):

no im learning this for the first time

OpenStudy (anonymous):

does it equal 0

OpenStudy (anonymous):

But here @vishweshshrimali5 you will use sin(A-B) formula..

OpenStudy (anonymous):

\[\sin(A-B) = \sin(A) \cdot \cos(B) - \cos(A) \cdot \sin(B)\] For 15 degree you will select A = 45 and B = 30..

OpenStudy (anonymous):

\[\sin(45 - 30) = \sin(45) \cdot \cos(30) - \cos(45) \cdot \sin(30)\]

OpenStudy (anonymous):

is it right to put sin 15deg =0 in simplest form

OpenStudy (anonymous):

or is it better left as is

OpenStudy (anonymous):

Now use this values and try to find out the value: \[\sin(45) = \frac{\sqrt{2}}{2}\] \[\cos(45) = \frac{\sqrt{2}}{2}\] \[\sin(30) = \frac{1}{2}\] \[\cos(30) = \frac{\sqrt{3}}{2}\] Put this values on Right Hand side and solve it.. Leave Left Hand Side as such..

OpenStudy (anonymous):

what value am i looking for is it pi/4, -pi/4, pi/6, -pi/6

OpenStudy (anonymous):

No.. Just do what I said.. Put those values in the formula that I gave you : \[\sin(45) \cdot \cos(30) - \cos(45) \cdot \sin(30)\] Here put those values..